OpenStudy (anonymous):

Evaluate 7 − 2y, if y = 5.

OpenStudy (anonymous):

what the question is asking is for you to plug 5 into the equation where y is. \[7-2(5)=7-10=-3\]

OpenStudy (anonymous):

does that answer your question?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

i got -3